OpenStudy (anonymous):

Write the polynomial function f(t) = - t - t(2t)^2 in standard form classify the function by degree and number of terms. A f(t) = 4t^2 - 2t; quadratic binomial b f(t) = 4t^2 - 2t; quadratic trinomial c f(t) = -4t^3 - t; cubic binomial d f(t) = - 4t^3 -t; cubic trinomial

OpenStudy (amoodarya):

-t-t*(2t)^2=-t -T*(2^2*t^2)= -t-t*(4*t^2)=-t-4*t^3 ok ?

OpenStudy (anonymous):

so is it D?

OpenStudy (anonymous):

@marsss Do you know this?

OpenStudy (anonymous):

i know that it's a cubic polynom but i don't know binomial or trinomial